A military aircraft is flying horizontally at a speed of 500m/s, drops a missile at an enemy target when it is at a height of 132.5m above the ground level.calculate in kilometers, the value of the horizontal distance travelled by the missile before hitting the target.[g= 10m/s square]


1
Expert's answer
2021-11-29T11:41:53-0500

Let's first find the time that the missile takes to reach the target:


"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\times132.5\ m}{10\ \dfrac{m}{s^2}}}=5.15\ s."

Finally, we can find the value of the horizontal distance travelled by the missile before hitting the target:


"x=v_0t=500\ \dfrac{m}{s}\times5.15\ s= 2575\ m."